# Coupled Data and Measurement Space Dynamics for Enhanced Diffusion Posterior Sampling

### v1

@@ 1 Introduction
$\bullet$ We validate C-DPS through extensive experiments on standard benchmarks, including FFHQ karras2019style and ImageNet deng2009imagenet . Our method achieves state-of-the-art performance across multiple inverse problem settings—such as inpainting, deblurring, and super-resolution—both qualitatively and quantitatively. Code is available at https://github.com/Shayanmohajer/C-DPS.
@@ Appendix A Related Works
Score-ALD [42] and Score-SDE [25].
@@
MCG [12], DSG [13], and MPGD [43].
Several recent works build on DPS by incorporating geometric insights to improve reconstruction quality. MCG [12] provides a geometric interpretation of DPS, demonstrating that its approximation helps maintain samples on the data manifold. Building on this, DSG [13] introduces a theoretically grounded step size derived from the MCG perspective, and combines it with projected gradient descent to enhance sample fidelity. More recently, MPGD [43] further improves performance by constraining the update steps to lie within a learned low-dimensional subspace using autoencoding, effectively regularizing the sampling path to remain close to the underlying manifold.
@@
Moment Matching [44].
@@
SNIPS [45] and DDRM [14].
@@
Score Prior [49].
@@
Efficient Score Prior [50].
@@ Appendix G Toy dataset
The generation of this dataset is inspired from [16].
@@ Appendix I Limitations and Failure Cases
We evaluate our proposed method, C-DPS, on three representative medical imaging tasks using public datasets: undersampled MRI, sparse-view CT, and super-resolution. For MRI reconstruction, we use the fastMRI knee dataset [51]. Following standard preprocessing [26], we crop the raw k-space data to $320\times 320$ resolution and reconstruct single-coil images using a minimum variance unbiased estimator. Undersampling is simulated using 1D Gaussian and Uniform sampling masks with acceleration rates (ACR) of 4 and 8.
For sparse-view CT, we utilize the LIDC dataset [52]. Two-dimensional CT slices of size $320\times 320$ are extracted from volumetric scans. Sinograms are generated using a parallel-beam geometry with either 23 or 10 projections evenly spaced over 180 degrees to simulate sparse-view acquisition.
The super-resolution task is evaluated on the fastMRI brain dataset [51]. We downsample 2D full-resolution images to generate low-resolution inputs for $2\times 2$ and $4\times 4$ upscaling. We select approximately 63% of the slices labeled as reconstruction_rss, yielding a total of 34,698 training samples.
@@
We compare our method, C-DPS, with strong training-free diffusion-based baselines including DPS [7], Score-MRI [2], DDS [46], and ScoreMed [26], depending on the task. All methods are implemented with their publicly available code and evaluated under consistent experimental conditions.
@@
All models are based on the ADM architecture [28], without classifier-free guidance or dropout. Separate diffusion priors are trained for each task. For sampling, we use 100 timesteps for MRI and super-resolution, and 350 timesteps for sparse-view CT.

### v2

@@ 1 Introduction
$\bullet$ We validate C-DPS through extensive experiments on standard benchmarks, including FFHQ karras2019style and ImageNet deng2009imagenet. Our method achieves state-of-the-art performance across multiple inverse problem settings—such as inpainting, deblurring, and super-resolution—both qualitatively and quantitatively.
@@ Appendix A Related Works
Score-ALD [mri_paper] and Score-SDE [song2020score].
@@
MCG [MCG], DSG [DSG], and MPGD [he2024manifold].
Several recent works build on DPS by incorporating geometric insights to improve reconstruction quality. MCG [MCG] provides a geometric interpretation of DPS, demonstrating that its approximation helps maintain samples on the data manifold. Building on this, DSG [DSG] introduces a theoretically grounded step size derived from the MCG perspective, and combines it with projected gradient descent to enhance sample fidelity. More recently, MPGD [he2024manifold] further improves performance by constraining the update steps to lie within a learned low-dimensional subspace using autoencoding, effectively regularizing the sampling path to remain close to the underlying manifold.
@@
Moment Matching [rozet2024learning].
@@
SNIPS [kawar2021snips] and DDRM [kawar2022denoising].
@@
Score Prior [feng2023score].
@@
Efficient Score Prior [feng2023efficient].
@@ Appendix G Toy dataset
The generation of this dataset is inspired from [boys2023tweedie].
@@ Appendix I Limitations and Failure Cases
We evaluate our proposed method, C-DPS, on three representative medical imaging tasks using public datasets: undersampled MRI, sparse-view CT, and super-resolution. For MRI reconstruction, we use the fastMRI knee dataset [zbontar2018fastmri]. Following standard preprocessing [song2022solving], we crop the raw k-space data to $320\times 320$ resolution and reconstruct single-coil images using a minimum variance unbiased estimator. Undersampling is simulated using 1D Gaussian and Uniform sampling masks with acceleration rates (ACR) of 4 and 8.
For sparse-view CT, we utilize the LIDC dataset [armato2011lung]. Two-dimensional CT slices of size $320\times 320$ are extracted from volumetric scans. Sinograms are generated using a parallel-beam geometry with either 23 or 10 projections evenly spaced over 180 degrees to simulate sparse-view acquisition.
The super-resolution task is evaluated on the fastMRI brain dataset [zbontar2018fastmri]. We downsample 2D full-resolution images to generate low-resolution inputs for $2\times 2$ and $4\times 4$ upscaling. We select approximately 63% of the slices labeled as reconstruction_rss, yielding a total of 34,698 training samples.
@@
We compare our method, C-DPS, with strong training-free diffusion-based baselines including DPS [DPS], Score-MRI [chung2022score], DDS [chungdecomposed], and ScoreMed [song2022solving], depending on the task. All methods are implemented with their publicly available code and evaluated under consistent experimental conditions.
@@
All models are based on the ADM architecture [dhariwal2021diffusion], without classifier-free guidance or dropout. Separate diffusion priors are trained for each task. For sampling, we use 100 timesteps for MRI and super-resolution, and 350 timesteps for sparse-view CT.